The problem is the following:
One number is four less than three times another. If their sum is increased by five, the result is twenty-five. Find the numbers.
I thought that maybe the equation would be x=3x-4 but it just doesn't work, then another way I thought would work would be x-4<3x but I don't know where to go from thier. I've been on this problem for an hour looking online for some help. I even went through a 40 minute free lesson online. It was too simple too help me though. Hopefully you can help me!

Answer by rfer(16322)

Give this a try.
(3x-4)+x=25-5
4x=24
x=6
3x-4=14